OpenStudy (anonymous):

i need help on this question: A path is determined by the equations x=1-(5/t^3) and y=2-(1/t^2) for t>0. For what time does the line tangent to this path have a slope of 8?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (nikvist):

\[x=1-5t^{-3}\quad\Rightarrow\quad dx=15t^{-4}\] \[y=2-t^{-2}\quad\Rightarrow\quad dy=2t^{-3}\] \[\frac{dy}{dx}=\frac{2t^{-3}}{15t^{-4}}=\frac{2t}{15}=8\quad\Rightarrow\quad t=60\]